# Theory of intervalley-coherent AFM order and topological superconductivity in tWSe2

### v1

@@ II.1 Multi-orbital Wannier model
where $\nu$ denotes the locked spin-valley degree of freedom, $X,X^{\prime}$ orbital, and $\bm{R},\bm{R}^{\prime}$ Bravais lattice vectors. The operator $c_{\bm{R}X}^{\nu\,(\dagger)}$ destroys (creates) an electron with spin/valley $\nu$ in the orbital $X$ on site $\bm{R}$, and $n^{\nu}_{\bm{R}X}=c^{\nu\,\dagger}_{\bm{R}X}c^{\nu{\vphantom{\dagger}}}_{\bm{R}X}$ is the density operator. As the hopping parameters $t_{\bm{R}X,\bm{R}^{\prime}X^{\prime}}^{\nu}$ are valley-diagonal, the interaction part is what couples different spins/valleys. To disentangle the microscopic mechanisms responsible for the formation of the Fermi surface reconstructed AFM phase and superconductivity observed in experimental measurements Xia et al. (2024); Guo et al. (2024), we first consider the dominant on-site interaction associated to the triangular ($U_{T}$) and honeycomb ($U_{H_{1,2}}$) sites and later investigate the influence of long-ranged density-density interaction on the phase diagram of $\theta=5.08^{\circ}$ tWSe2. The values of $U_{T}\approx 74\,\mathrm{meV}$ and $U_{H}\approx 106\,\mathrm{meV}$ place tWSe2 into the weak-to-intermediate coupling regime, where the band width $W\approx 200\,\mathrm{meV}$ of the three-orbital model fulfills $W\gtrsim U_{T/H}$. We note that due to the topological entanglement of the topmost valence bands in tWSe2, the bandwidth of the entire Wannier band manifold (and not only the topmost valence band) should be considered.
Figure 2: FRG phase diagram of $\theta=5.08^{\circ}$ tWSe2 resolving the interplay of IVC-AFM order and chiral $d/p$-wave superconductivity along the gate-tunable van-Hove singularity. (a) Main panel: FRG phase diagram as function of holes per moiré unit cell $n/n_{0}$ and external displacement field $E_{z}$. The phase diagram shows the critical scale $\Lambda_{\mathrm{c}}$ of the leading Fermi surface instability and indicates regimes of spin/charge density wave order (DW, red) and superconductivity (SC, blue). The density of states (DOS) is shown in the background (gray) to pinpoint the position of the van-Hove line. Close to the HOVHS (blue circle), the system features IVC-AFM order, while for lower values of the displacement field chiral $d/p$-wave superconductivity emerges on the van-Hove line. Inset: Zoom-in on the density-displacement field region where IVC-AFM order gives way to superconductivity. The colored arrows indicate different directions when moving along the van-Hove line (green arrows) or at constant displacement field cuts (grey, blue arrows). The critical scale $\Lambda_{\mathrm{c}}$ along these line cuts is shown in the panels on the left, indicating a maximal critical temperature of $\Lambda_{\mathrm{c}}\sim 500\,\mathrm{mK}$ that is reached in the immediate vicinity of the IVC-AFM phase. (b) Momentum structure and (in)commensurability of the IVC-AFM state along the van-Hove line. The subpanels show the absolute value of the particle-hole susceptibility $\chi^{\text{FRG}}_{\uparrow\downarrow\uparrow\downarrow}(\bm{q})$ at scale $\Lambda_{\mathrm{c}}$ for the relevant orbital components involving the Wannier orbitals $T/H_{2}$. The leading momentum transfer $\bm{Q}_{C}$ locks to $K^{\prime}$ at the HOVHS (blue circle) and successively shifts towards incommensurate momenta along the high symmetry path $K^{\prime}$-$M$ when lowering the external displacement field along the van-Hove line (blue square). (c) Amplitude and phase dependence of the pairing instability at $E_{z}=25\,\mathrm{meV}$ (blue circle) decomposed into singlet $\psi_{0}(\bm{k})$ and triplet $d_{z}(\bm{k})$ components. The continuous colorbar indicates the relative phase, whereas the amplitude is encoded by the opacity. The SC order parameter is two-fold degenerate and likely minimizes its free energy by the chiral superposition $d+id/p+ip$.
II.2 Functional renormalization group approach
In the weak-to-moderate coupling regime, the functional renormalization group (FRG) Salmhofer and Honerkamp (2001); Metzner et al. (2012); Platt et al. (2013); Dupuis et al. (2021); Beyer et al. (2022) represents a well-established method to predict particle-particle and particle-hole instabilities in an unbiased manner. FRG smoothly connects the non-interacting action to the interacting one by virtue of a renormalization group (RG) scale $\Lambda=\infty\to 0$. The introduction of the RG scale leads to an infinite hierarchy of coupled differential equations for the vertex functions $\Gamma^{(2n)}$. In order to numerically solve the coupled nonlinear functional differential equations, we must truncate the hierarchy and approximate the vertex functions’ dependencies on frequency and momentum. To account for the topology of the multi-orbital Wannier setup for the moiré valence bands, we here employ the static four-point truncated unity FRG (TUFRG) approximation Husemann and Salmhofer (2009); Lichtenstein et al. (2017); Profe and Kennes (2022) that allows to keep track of the entire orbital/band and (incommensurate) momentum dependence of the model at the expense of neglecting the frequency dependence of the vertex functions. Most importantly, this allows to treat all couplings on equal footing and (at one loop order) accounts for mutual feedback between the different interaction channels that are renormalized during the flow. The vertex functions are parametrized in a hybrid real-momentum space representation Profe and Kennes (2022); Profe et al. (2024) using a $30\times 30$ momentum mesh for the bosonic momentum variable in each two-particle reducible interaction channel—hence allowing to resolve (in)commensurate particle-particle and particle-hole instabilities—whereas the fermionic momenta are expanded in a real-space form factor basis that is truncated after a distance of 2$|\bm{a}_{M}|$. The propagator is resolved on a refined $900\times 900$ momentum mesh to increase the energy resolution and to sufficiently resolve the fermiology of the system at low scales, see SM SM for details. In static four-point (TU)FRG, the onset of an ordered phase is signaled by the divergence of the two-particle vertex $\Gamma^{(4)}$ at a particular (critical) scale $\Lambda_{\mathrm{c}}$. The most divergent eigenvalue (and the accompanying eigenvector) of $\Gamma^{(4)}$ then provides insights into the structure of the ordered phase—whether it is a particle-particle or a particle-hole instability as well as the momentum, spin/valley and orbital structure of the order parameter. Since we choose the sharp frequency cutoff as regulator (as implemented in the divERGe library Profe et al. (2024)), the critical scale $\Lambda_{c}$ serves as a proxy for the critical temperature of the transition (further details on FRG are given in the SM SM ).
III FRG phase diagram
We present the FRG phase diagram of $5.08^{\circ}$ tWSe2 in Fig. 2 (a), where we plot the critical scale $\Lambda_{\mathrm{c}}$ of emergent particle-particle (blue) and particle-hole (red) instabilities as a function of hole filling per moiré unit cell $n/n_{0}$ and transverse electrical field $E_{z}$. The density of states (DOS) is shown as a reference to pinpoint the position of the van-Hove line (gray). At the position of the HOVHS, the system promotes an inter-valley coherent (IVC) anti-ferromagnetic spin-density wave with order parameter
@@ III FRG phase diagram
where $\vec{\sigma}$ acts in the space of the locked spin-valley degree of freedom and $\vartheta_{X}$ measures the relative orientation and weight of the $120^{\circ}$ in-plane spin orientation on the orbitals $T/H_{2}$ that is otherwise arbitrary due to the remaining $U_{\nu}(1)$ symmetry. The spin density wave breaks translational invariance due to the finite transfer momentum $\bm{Q}_{C}=K^{\nu}$ that connects the HOVHSs in different spin-valley sectors and transforms in the trivial $A$ irreducible representation (IR) of the $C_{3z}$ little group at $K^{\nu}$. Consistent with the orbital polarization of the Fermi pockets at the HOVHS, the order parameter has dominant weight on the $T/H_{2}$-orbitals as imprinted in the transverse particle-hole susceptibility $\chi_{\uparrow\downarrow\uparrow\downarrow}^{\text{FRG}}(\bm{q})$ (see SM SM for details) shown in Fig. 2 (b). Decreasing the value of the displacement field $E_{z}$ along the van-Hove line shifts the leading momentum transfer $\bm{Q}_{C}$ towards the $M$-point, causing regions of incommensurate IVC-AFM order as demonstrated in Fig. 2 (b).
@@
The absence of a perfect scattering vector that connects the van-Hove points in different spin-valley sectors and the reduction of the DOS weakens particle-hole instabilities for $E_{z}<E_{z}^{c}$ and instead promotes the effect of particle-particle fluctuations leading to the formation of superconductivity along the van-Hove line, see blue square in Fig. 2 (a). The formation of superconductivity along the van-Hove line is hence facilitated by Fermi surface warping when detuning the system from the HOVHS, reminiscent of pairing in the $t$-$t^{\prime}$ square lattice Hubbard model Honerkamp (2003) as will be discussed in Section IV.1 in more detail.
The Ising-like spin-orbit coupling in tWSe2 couples singlet- and triplet components of the superconducting order parameter
@@
where the hat-symbol encodes the orbital and spin-valley degree of freedom. We find that the leading superconducting order parameter is an inter-valley symmetric combination ($S^{z}=0$) of singlet and triplet components $[\psi_{0}(\bm{k}),d_{z}(\bm{k})]$ that are allowed to admix due to the preserved $U_{\nu}(1)$ symmetry Schrade and Fu (2021). The Fermi surface projected amplitude and phase of the respective components is shown in the right panel of Fig. 2 (c). The leading order parameter transforms in the two-dimensional $E$ irreducible representation of the Hamiltonian’s point group $\mathcal{G}_{0}$. In a mean-field picture, the two degenerate order parameters are in general superimposed and the free energy is presumably minimized by the chiral, time-reversal breaking combination $d\pm id/p\pm ip$. The latter gaps the Fermi surface and hence minimizes the energy of the superconducting condensate Honerkamp (2003); Nandkishore et al. (2012); Black-Schaffer and Honerkamp (2014); Classen et al. (2019); Scherer et al. (2022). The chiral superconducting state belongs to the A Cartan (Altland-Zirnbauer) class for $S_{z}$ preserving Bogoliubov de-Gennes Hamiltonians Schnyder et al. (2008) possessing a $\mathbb{Z}$ topological invariant that manifests in a non-vanishing Chern number $C=\pm 2$ and edge modes that are described by Dirac-like complex fermionic quasiparticle.

### v2

@@ III.1 Multi-orbital Wannier model
III.2 Dual-gated Coulomb interactions
To model the long-ranged Coulomb interaction between electrons in tWSe2, we employ an experimentally motivated interaction profile $V(r)$ that accounts for the dual-gated device architecture as well as short-ranged interactions that are regularized by an Ohno form [86, 79]:
$V(r)=4V_{0}\sum_{k=0}^{\infty}K_{0}\left[(2k+1)\pi\,\frac{\sqrt{r^{2}+a^{2}}}{\xi}\right]\,.$ (1)
Here, $K_{0}$ denotes a modified Bessel function of the second kind, $\xi$ is the distance from the sample to either gate and $\alpha=14.40$ eV Å is the fine-structure constant. The interaction parameters $a=\alpha/(\epsilon U)$ and $V_{0}=\alpha/(\epsilon\xi)$, only depend on the dielectric constant $\epsilon$ and the on-site interaction strength $U$, see SM [76] for details. Motivated by the experimental gating architecture [37, 39], we choose $\xi=100$ Å and treat the exact value of the dielectric constant as parameter as it further depends on internal screening in tWSe2 mediated by remote bands and external screening by the hBN encapsulating layers. In the main text, we choose $\epsilon=16$ in agreement with first-principle estimates [87], however, provide additional data for reasonable values of $\epsilon$ in the SM [76] that demonstrate the robustness of our results.
The dual-gated Coulomb interaction defined in Eq. (1) is then projected onto the multi-orbital Wannier model, giving rise to an effective low-energy Hamiltonian of the form
@@ III.2 Dual-gated Coulomb interactions
where $\nu$ denotes the locked spin-valley degree of freedom, $X,X^{\prime}$ orbital, and $\bm{R},\bm{R}^{\prime}$ Bravais lattice vectors. The operator $c_{\bm{R}X}^{\nu\,(\dagger)}$ destroys (creates) an electron with spin/valley $\nu$ in the orbital $X$ on site $\bm{R}$, and $n^{\nu}_{\bm{R}X}=c^{\nu\,\dagger}_{\bm{R}X}c^{\nu{\vphantom{\dagger}}}_{\bm{R}X}$ is the density operator. As the hopping parameters $t_{\bm{R}X,\bm{R}^{\prime}X^{\prime}}^{\nu}$ are valley-diagonal, the interaction part is what couples different spins/valleys. To disentangle the microscopic mechanisms responsible for the formation of the Fermi surface reconstructed AFM phase and superconductivity observed in experimental measurements [37, 39], we first consider the dominant on-site interaction associated to the triangular ($U_{T}$) and honeycomb ($U_{H_{1,2}}$) sites and later investigate the influence of long-ranged density-density interaction on the phase diagram of $\theta=5.08^{\circ}$ tWSe2. We further provide a detailed list of the effective interaction parameters in the Wannier model for different values of $(\xi,\epsilon)$ in the SM [76]. The values of $U_{T}\approx 74\,\mathrm{meV}$ and $U_{H}\approx 106\,\mathrm{meV}$ place tWSe2 into the weak-to-intermediate coupling regime, where the band width $W\approx 200\,\mathrm{meV}$ of the three-orbital model fulfills $W\gtrsim U_{T/H}$. We note that due to the topological entanglement of the topmost valence bands in tWSe2, the bandwidth of the entire Wannier band manifold (and not only the topmost valence band) should be considered.
Figure 2: FRG phase diagram of $\theta=5.08^{\circ}$ tWSe2 resolving the interplay of IVC-AFM order and chiral $d/p$-wave superconductivity along the gate-tunable van-Hove singularity. (a) Main panel: FRG phase diagram as function of holes per moiré unit cell $n/n_{0}$ and external displacement field $E_{z}$. The phase diagram shows the critical scale $\Lambda_{\mathrm{c}}$ of the leading Fermi surface instability and indicates regimes of spin/charge density wave order (DW, red) and superconductivity (SC, blue). The density of states (DOS) is shown in the background (gray) to pinpoint the position of the van-Hove line. Close to the HOVHS (blue circle), the system features IVC-AFM order, while for lower values of the displacement field chiral $d/p$-wave superconductivity emerges on the van-Hove line. Inset: Zoom-in on the density-displacement field region where IVC-AFM order gives way to superconductivity. The colored arrows indicate different directions when moving along the van-Hove line (green arrows) or at constant displacement field cuts (grey, blue arrows). The critical scale $\Lambda_{\mathrm{c}}$ along these line cuts is shown in the panels on the left, indicating a maximal critical temperature of $\Lambda_{\mathrm{c}}\sim 500\,\mathrm{mK}$ that is reached in the immediate vicinity of the IVC-AFM phase. (b) Momentum structure and (in)commensurability of the IVC-AFM state along the van-Hove line. The subpanels show the absolute value of the inter-valley particle-hole susceptibility $\chi^{\text{FRG}}_{\uparrow\downarrow\uparrow\downarrow}(\bm{q})$ at scale $\Lambda_{\mathrm{c}}$ for the relevant orbital components involving the Wannier orbitals $T/H_{2}$. The leading momentum transfer $\bm{Q}_{C}$ locks to $K^{\prime}$ at the HOVHS (blue circle) and successively shifts towards incommensurate momenta along the high symmetry path $K^{\prime}$-$M$ when lowering the external displacement field along the van-Hove line (blue square). (c) Amplitude and phase dependence of the pairing instability at $E_{z}=25\,\mathrm{meV}$ (blue circle) decomposed into singlet $\psi_{0}(\bm{k})$ and triplet $d_{z}(\bm{k})$ components. The continuous colorbar indicates the relative phase, whereas the amplitude is encoded by the opacity. The SC order parameter is two-fold degenerate and likely minimizes its free energy by the chiral superposition $d+id/p+ip$.
III.3 Functional renormalization group approach
In the weak-to-moderate coupling regime, the functional renormalization group (FRG) [60, 61, 62, 63, 64] represents a well-established method to predict particle-particle and particle-hole instabilities in an unbiased manner. FRG smoothly connects the non-interacting action to the interacting one by virtue of a renormalization group (RG) scale $\Lambda=\infty\to 0$. The introduction of the RG scale leads to an infinite hierarchy of coupled differential equations for the vertex functions $\Gamma^{(2n)}$. In order to numerically solve the coupled nonlinear differential equations, we must truncate the hierarchy and approximate the vertex functions’ dependencies on frequency and momentum. To account for the topology of the multi-orbital Wannier setup for the moiré valence bands, we here employ the static four-point truncated unity FRG (TUFRG) approximation [88, 89, 90] that allows to keep track of the entire orbital/band and (incommensurate) momentum dependence of the model at the expense of neglecting the frequency dependence of the vertex functions. The level-2 truncated FRG flow equations for the two-particle vertex function $\Gamma=\Gamma^{(4)}$ read
$\partial_{\Lambda}\Gamma^{\Lambda}=\sum_{\gamma}\partial_{\Lambda}\gamma^{\Lambda}\,,\quad\partial_{\Lambda}\gamma^{\Lambda}=\Gamma^{\Lambda}\circ\partial_{\Lambda}L^{\gamma,\Lambda}\circ\Gamma^{\Lambda}\,$, (3)
where $\gamma\in\{P,D,C\}$ denotes the decomposition into two-particle reducible interaction channels that capture fluctuations accumulated in the particle-particle ($P$), direct particle-hole ($D$) and crossed particle-hole ($C$) channel. Most importantly, this allows to treat all couplings on equal footing and (at one loop order) accounts for mutual feedback between the different interaction channels that are renormalized during the flow. In static four-point (TU)FRG, the onset of an ordered phase is signaled by the divergence of the two-particle vertex $\Gamma^{(4)}$ at a particular (critical) scale $\Lambda_{\mathrm{c}}$. The leading fermionic bilinear in the divergent channel is obtained by an eigenvalue decomposition of the channel-specific vertex $\gamma$
$\gamma_{\kappa\kappa^{\prime}}(\bm{q}_{\gamma})=\sum_{i}\phi^{L}_{\kappa,i}(\bm{q}_{\gamma})\,\gamma_{i}(\bm{q}_{\gamma})\,\phi^{R}_{\kappa^{\prime},i}(\bm{q}_{\gamma})\,$, (4)
where $\kappa=(\bm{k}_{\gamma},X,\nu_{1},\nu_{2})$ is a multi-index comprising the fermionic momentum variable $\bm{k}_{\gamma}$, Wannier function index $X$, and spin/valley $\nu_{1,2}$. The left/right eigenvectors to the channels’ eigenvalues $\gamma_{i}$ are denoted by $\phi^{L/R}_{\kappa,i}$. They provide insight into the momentum, spin/valley and orbital structure of the order parameter. In the absence of spin $SU(2)$-symmetry, particle-hole instabilities in the $C/D$-channel are allowed to admix and we therefore refer to them as density wave instabilities (DW) in the following. Since we choose the sharp frequency cutoff as regulator (as implemented in the divERGe library [91]), the critical scale $\Lambda_{c}$ serves as a proxy for the critical temperature of the transition. To render the solution of Eq. (3) numerically feasible, the vertex functions are parametrized in a hybrid real-momentum space representation [90, 91] using a $30\times 30$ momentum mesh for the bosonic momentum variable ($\bm{q}_{\gamma}$) in each two-particle reducible interaction channel—hence allowing to resolve (in)commensurate particle-particle and particle-hole instabilities—whereas the fermionic momenta ($\bm{k}_{\gamma}$) are expanded in a real-space form factor basis that is truncated after a distance of $2\ldots 4|\bm{a}_{M}|$. The two-particle propagator $L^{\gamma,\Lambda}$ is resolved on a refined $900\times 900$ momentum mesh to increase the energy resolution and to sufficiently resolve the fermiology of the system at low scales. A detailed overview of the level-2 truncated FRG flow equations and technical details are provided in the SM [76].
The level-2 truncation of the FRG flow adapted in Eq. (3) breaks down when approaching the symmetry-broken phase as indicated by a divergence of the two-particle vertex $\Gamma^{\Lambda}$ at scale $\Lambda=\Lambda_{\mathrm{c}}$. To access quasiparticle properties and spectral functions in the ordered phase for scales $\Lambda_{\text{MF}}\lesssim\Lambda_{\mathrm{c}}$ predicted by FRG, we may formulate an effective mean-field (MF) theory as outlined in Ref. [61, 92]. To overcome the non-analytic behavior of the FRG flow at the critical scale $\Lambda_{\mathrm{c}}$ and connect the symmetry-broken state with the high-temperature normal-state, we construct a suitable bare interaction $V^{\text{MF}}$ for MF treatment by solving an inverse Bethe-Salpeter equation in the divergent channel at scale $\Lambda_{\mathrm{c}}$
$\gamma=V^{\text{MF}}+V^{\text{MF}}\circ\chi^{0,\gamma}\circ\gamma\,$, (5)
where $\chi^{0,\gamma}$ is the bare particle-particle (particle-hole) susceptibility, i.e $\chi^{0,\gamma}=\int^{\infty}_{\Lambda_{\mathrm{c}}}\text{d}\Lambda\,\partial_{\Lambda}L^{\gamma,\Lambda}$. Inverting this equation yields the bare interaction $V^{\text{MF}}$ that by construction yields the correct MF state upon a successive self-consistent MF decoupling. We further note that the single-channel RPA approximation can be recovered from the full FRG flow equation defined in Eq. (3) by neglecting inter-channel coupling, i.e. only considering the single-channel ladder resummation and integrating the second term in Eq. (3)
$\gamma^{\text{RPA}}=\Gamma^{0}\circ\chi^{0,\gamma}\circ\Gamma^{0}\,$, (6)
where $\Gamma^{0}$ refers to the initial (bare) interaction.
IV FRG phase diagram
We present the FRG phase diagram of $5.08^{\circ}$ tWSe2 in Fig.˜2 (a), where we plot the critical scale $\Lambda_{\mathrm{c}}$ of emergent particle-particle (SC, blue) and particle-hole (DW, red) instabilities as a function of hole filling per moiré unit cell $n/n_{0}$ and transverse electrical field $E_{z}$. The density of states (DOS) is shown as a reference to pinpoint the position of the van-Hove line (gray). At the position of the HOVHS, the system promotes an inter-valley coherent (IVC) anti-ferromagnetic spin-density wave with order parameter
$\vec{\Delta}^{\text{AFM}}_{\bm{R}}\propto\!\!\!\sum_{X\in\{T,H_{2}\}}e^{i[\bm{Q}_{C}\cdot\bm{R}+i\vartheta_{X}]}\langle c^{\nu\,\dagger}_{\bm{R}X}\vec{\sigma}^{\nu\nu^{\prime}}c^{\nu^{\prime}{\vphantom{\dagger}}}_{\bm{R}X}\rangle+\text{h.c.}\,$, (7)
@@ IV FRG phase diagram
where $\vec{\sigma}=(\sigma_{x},\sigma_{y})^{\mathrm{T}}$ acts in the space of the locked spin-valley degree of freedom and $\vartheta_{X}$ measures the relative orientation and weight of the $120^{\circ}$ in-plane spin orientation on the orbitals $T/H_{2}$ that is otherwise arbitrary due to the remaining $U_{\nu}(1)$ symmetry. The spin density wave ($\mathcal{T}$-odd) breaks translational invariance due to the finite transfer momentum $\bm{Q}_{C}=K^{\nu}$ that connects the HOVHSs in different spin-valley sectors and transforms in the trivial $A$ irreducible representation (IR) of the $C_{3z}$ little group at $K^{\nu}$. Consistent with the orbital polarization of the Fermi pockets at the HOVHS, the order parameter has dominant weight on the $T/H_{2}$-orbitals as imprinted in the inter-valley particle-hole susceptibility $\chi_{\uparrow\downarrow\uparrow\downarrow}^{\text{FRG}}(\bm{q})$ (see SM [76] for details) shown in Fig.˜2 (b). We further provide a sketch of the real-space spin pattern in Fig. 4 (a). Decreasing the value of the displacement field $E_{z}$ along the van-Hove line shifts the leading momentum transfer $\bm{Q}_{C}$ towards the $M$-point, causing regions of incommensurate IVC-AFM order as demonstrated in Fig.˜2 (b).
@@
The absence of a perfect scattering vector that connects the van-Hove points in different spin-valley sectors and the reduction of the DOS weakens particle-hole instabilities for $E_{z}<E_{z}^{c}$ and instead promotes the effect of particle-particle fluctuations leading to the formation of superconductivity along the van-Hove line, see blue square in Fig.˜2 (a). The formation of superconductivity along the van-Hove line is hence facilitated by Fermi surface warping when detuning the system from the HOVHS, reminiscent of pairing in the $tt^{\prime}$ square lattice Hubbard model [93] as will be discussed in Section˜VI.1 in more detail. The Ising-like spin-orbit coupling in tWSe2 couples singlet- and triplet components of the superconducting order parameter
@@
where the hat-symbol encodes the orbital and spin-valley degree of freedom. We find that the leading superconducting order parameter is an inter-valley symmetric combination ($S^{z}=0$) of singlet and triplet components $\psi_{0}(\bm{k}),d_{z}(\bm{k})]$ that are allowed to admix due to the preserved $U_{\nu}(1)$ symmetry [40]. The Fermi surface projected amplitude and phase of the respective components is shown in the right panel of Fig.˜2 (c). Similar to the IVC-AFM order parameter, the SC order parameter only has weight on bonds including the $T/H_{2}$-orbital, whereas $\Delta_{H_{1}X}(\bm{k})=\Delta_{XH_{1}}(\bm{k})=0$ for $X\in\{T,H_{1},H_{2}\}$, see SM [76] for full orbital-resolved structure of the SC order parameter. The leading order parameter transforms in the two-dimensional $E$ irreducible representation of the Hamiltonian’s point group $\mathcal{G}_{0}$. In a mean-field picture, the two degenerate order parameters are in general superimposed and the free energy is minimized by the chiral, time-reversal breaking combination $d\pm id/p\pm ip$. The latter gaps the Fermi surface and hence minimizes the energy of the superconducting condensate [93, 94, 95, 96, 97]. The chiral superconducting state belongs to the A Cartan (Altland-Zirnbauer) class for $S_{z}$ preserving Bogoliubov de-Gennes Hamiltonians [98] possessing a $\mathbb{Z}$ topological invariant that manifests in a non-vanishing Chern number $C=\pm 2$ and edge modes that are described by Dirac-like complex fermionic quasiparticle.